fined and equation of the line that satifies the given conditions through (1,7) parallel to the line passing through (2,5) and (-2,1)?
Answer by ewatrrr(24785)

parallel to the line passing through
(2,5) and
(-2,1) m = 4 / 4 = 1
New parallel Line passes thru(1,7)
y - 7 = 1(x-1)
y = x + 6 0r x - y = -6 in Standard Form